Supposex ~N(0,1) and Y=X2 Write down the p.d.f. of Y. Calculate E(Y) and Var(Y).


1
Expert's answer
2021-06-29T07:59:55-0400

If the random variable "V \sim N(\mu, \sigma^2),\sigma^2>0," then the random variable "W=\dfrac{(X-\mu)^2}{\sigma^2}=Z^2\sim \chi^2(1)," chi-squared distribution, with "\nu=1" degrees of freedom.


Suppose "X\sim N(0, 1)," then "Y=\dfrac{(X-0)^2}{1^2}=X^2\sim \chi^2(1)," chi-squared distribution, with "\nu=1" degrees of freedom.

Its density function is given by


"f(y;\nu)= \begin{cases}
 \dfrac{1}{2^{\nu/2}\Gamma(\nu/2)}y^{\nu/2-1}e^{-y/2},y>0 \\
 0, \text{elsewhere} 
\end{cases}"

Then


"f(y;\nu)= \begin{cases}
 \dfrac{1}{\sqrt{2\pi}\sqrt{y}}e^{-y/2},y>0 \\
 0, \text{elsewhere} 
\end{cases}"

The mean of the distribution is equal to the number of degrees of freedom:


"E(Y)=\mu_Y=\nu=1"

The variance is equal to two times the number of degrees of freedom:


"Var(Y)=\sigma_Y^2=2\nu=2"